# Infinitesimal Variations of Hodge Structure for Singular Curves I

### v1

@@ 1. Introduction
In the classical smooth setting, maximality of the IVHS is expected and often holds generically. For example, for a general curve of genus $g$, the associated infinitesimal variation of Hodge structure has maximal rank $g^{2}$, reflecting the fact that the period map is generically immersive. This phenomenon underlies many deep results in the geometry of moduli spaces and plays a key role in the study of algebraic cycles and Hodge loci.
@@
Recently, in [14] Sernesi has obtained important results on the infinitesimal variation of Hodge structure for families of nodal plane curves, establishing generic maximality under suitable numerical conditions through a global deformation-theoretic and Hodge-theoretic analysis. The present work should be viewed as complementary to Sernesi’s contribution. While both works address closely related questions concerning maximal variation, they rely on fundamentally different approaches. In contrast with Sernesi’s global methods, our analysis is based on a local and residue-theoretic description of the IVHS, leading to a canonical decomposition of the variation map into contributions supported at singular points
@@
    If all singularities are nodes and cusps and the number of nodes is at least $g$, then $\Phi$ has maximal rank $g^{2}$.
@@
has maximal rank $g^{2}$.
@@
    If all singularities of $\mathcal{C}$ are nodes and cusps, and if the number of nodes satisfies $\delta\geq g$, then the infinitesimal variation of Hodge structure associated with the normalization family $\varphi\colon\mathcal{C}\longrightarrow\Delta$ has maximal variation. More precisely,
    $\operatorname{rank}(\Phi_{\varphi})=g^{2}.$
@@ 2. Preliminaries
Dualizing sheaf and normalization. Let $\omega_{\mathcal{C}}$ denote the dualizing sheaf of the possibly singular curve $\mathcal{C}$. By definition, $\omega_{\mathcal{C}}$ represents Serre duality (for more details see [10], [11], and [12]):
@@ 2.1. Explicit local models: nodes and cusps
summed over all $p\in{\operatorname{Sing}}(\mathcal{C})$ (see [9], [7]).
@@ 3.3. Local decomposition
    If all singularities are nodes and cusps and the number of nodes is at least $g$, then $\Phi$ has maximal rank $g^{2}$.
@@ 4.4. Local IVHS contributions
11. Local analytic model of the cusp. Since $Q_{j}$ is a cusp, there exist analytic coordinates $(x,y)$ centered at $Q_{j}$ such that $\mathcal{C}$ is locally given by $y^{2}=x^{3}.$ The normalization map $\varphi:C\to\mathcal{C}$ is given locally by the ramified parametrization $x=t^{2},$   $y=t^{3},$ where $t$ is a local coordinate on $C$ centered at $r_{j}$ (cf [1]).
@@
where $\xi_{P_{i}}$ denotes the component of $\xi$ corresponding to smoothing (or infinitesimal deformation) at the node $P_{i}$ (see for example [3], and [6]).
@@ 4.5. Residues and Cayley–Bacharach
Thus, a holomorphic differential on $C$ corresponds to a plane curve of degree d3d-3 passing through all nodes of $\mathcal{C}$.
@@
44. Cayley–Bacharach for nodal schemes. The scheme $\Delta$ of nodes of an irreducible plane curve satisfies the Cayley–Bacharach property with respect to curves of degree d3d-3. Concretely, this means that any plane curve of degree d3d-3 which vanishes to order at least two at all but one node must vanish to order at least two at the remaining node as well (we can see or example [5], [2], and [4]).
@@ 5.1. Maximal–variation theorem
    If all singularities of $\mathcal{C}$ are nodes and cusps, and if the number of nodes satisfies $\delta\geq g$, then the infinitesimal variation of Hodge structure associated with the normalization family $\varphi\colon\mathcal{C}\longrightarrow\Delta$ has maximal variation. More precisely,
    $\operatorname{rank}(\Phi_{\varphi})=g^{2}.$
@@ 5.2. Adjoint linear systems and double vanishing at nodes
consisting of adjoint sections vanishing to order at least two at every node ( see for example [13]).
@@ 5.3. Statement of the Cayley–Bacharach property
The Cayley–Bacharach property for the nodal scheme $\Delta$ with respect to the linear system $|\omega_{S}\otimes\mathcal{O}_{S}(\mathcal{C})|$ asserts the following   (see [2]):
@@ 5.6. Conclusion
As a consequence, holomorphic differentials on $C$ are identified with sections of the adjoint line bundle $\omega_{S}\otimes\mathcal{O}_{S}(\mathcal{C})$ whose restrictions to $\mathcal{C}$ satisfy the local compatibility conditions at the singular points (see for example [8], or [10]).
@@ 5.8. Failure of surjectivity for ADE singularities
$\operatorname{rank}(\Phi_{\varphi})\leq\delta<g^{2}$
for $g$ sufficiently large, even if $\delta$ grows linearly with $g$. Thus $\Phi_{\varphi}$ is not surjective, and maximal variation fails.

### v2

@@ 1. Introduction
In the classical smooth setting, maximality of the IVHS is expected and often holds generically. For example, for a general curve of genus $g$, the associated infinitesimal variation of Hodge structure has maximal rank, reflecting the fact that the period map is generically immersive. This phenomenon underlies many deep results in the geometry of moduli spaces and plays a key role in the study of algebraic cycles and Hodge loci.
@@
Recently, Sernesi has obtained important results on the infinitesimal variation of Hodge structure for families of nodal plane curves, establishing generic maximality under suitable numerical conditions through a global deformation-theoretic and Hodge-theoretic analysis. The present work should be viewed as complementary to Sernesi’s contribution. While both works address closely related questions concerning maximal variation, they rely on fundamentally different approaches. In contrast with Sernesi’s global methods, our analysis is based on a local and residue-theoretic description of the IVHS, leading to a canonical decomposition of the variation map into contributions supported at singular points
@@
    If all singularities are nodes and cusps and the number of nodes is at least $g$, then $\Phi$ has maximal rank (i.e. $\dim H^{0}(C,N_{\varphi})$).
@@
is injective (has maximal rank equal $\dim H^{0}(C,N_{\varphi})$).
@@
    If all singularities of $\mathcal{C}$ are nodes and cusps, and if the number of nodes satisfies $\delta\geq g$, then the infinitesimal variation of Hodge structure associated with the normalization family $\varphi\colon\mathcal{C}\longrightarrow\Delta$ has maximal variation (i.e. injective).
@@ 2. Preliminaries
Dualizing sheaf and normalization. Let $\omega_{\mathcal{C}}$ denote the dualizing sheaf of the possibly singular curve $\mathcal{C}$. By definition, $\omega_{\mathcal{C}}$ represents Serre duality:
@@ 2.1. Explicit local models: nodes and cusps
summed over all $p\in{\operatorname{Sing}}(\mathcal{C})$.
@@ 3.3. Local decomposition
    If all singularities are nodes and cusps and the number of nodes is at least $g$, then $\Phi$ has maximal rank.
@@ 4.4. Local IVHS contributions
11. Local analytic model of the cusp. Since $Q_{j}$ is a cusp, there exist analytic coordinates $(x,y)$ centered at $Q_{j}$ such that $\mathcal{C}$ is locally given by $y^{2}=x^{3}.$ The normalization map $\varphi:C\to\mathcal{C}$ is given locally by the ramified parametrization $x=t^{2},$   $y=t^{3},$ where $t$ is a local coordinate on $C$ centered at $r_{j}$.
@@
where $\xi_{P_{i}}$ denotes the component of $\xi$ corresponding to smoothing (or infinitesimal deformation) at the node $P_{i}$.
@@ 4.5. Residues and Cayley–Bacharach
Thus a holomorphic differential on $C$ corresponds to a plane curve of degree d3d-3 passing through all nodes of $\mathcal{C}$.
@@
44. Cayley–Bacharach for nodal schemes. The scheme $\Delta$ of nodes of an irreducible plane curve satisfies the Cayley–Bacharach property with respect to curves of degree d3d-3. Concretely, this means that any plane curve of degree d3d-3 which vanishes to order at least two at all but one node must vanish to order at least two at the remaining node as well.
@@ 5.1. Maximal–variation theorem
    If all singularities of $\mathcal{C}$ are nodes and cusps, and if the number of nodes satisfies $\delta\geq g$, then the infinitesimal variation of Hodge structure associated with the normalization family $\varphi\colon\mathcal{C}\longrightarrow\Delta$ has maximal variation.
@@ 5.2. Adjoint linear systems and double vanishing at nodes
consisting of adjoint sections vanishing to order at least two at every node.
@@ 5.3. Statement of the Cayley–Bacharach property
The Cayley–Bacharach property for the nodal scheme $\Delta$ with respect to the linear system $|\omega_{S}\otimes\mathcal{O}_{S}(\mathcal{C})|$ asserts the following   (see [7] or [3]):
@@ 5.6. Conclusion
As a consequence, holomorphic differentials on $C$ are identified with sections of the adjoint line bundle $\omega_{S}\otimes\mathcal{O}_{S}(\mathcal{C})$ whose restrictions to $\mathcal{C}$ satisfy the local compatibility conditions at the singular points.
@@ 5.8. Failure of surjectivity for ADE singularities
$\operatorname{rank}(\Phi_{\varphi})\leq\delta<3d-1+g$
for $g$ sufficiently large, even if $\delta$ grows linearly with $g$. Thus $\Phi_{\varphi}$ is not injective, and maximal variation fails.